OpenStudy (madeleine2):

What must be added to 3x -7 to make x^ 2 + 4x -1?

rishavraj (rishavraj):

3x - 7 + (y) = x^2 + 4x - 1 y = x^2 + 4x - 1 - 3x + 7 find y...thts ur answer :)